OpenStudy (anonymous):

sin theta _____ = 1/2 [sin(a+b)+sin(a-b)] A. cos b B. cos a C. sin a D. sin b

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

i'm not sure you typed that out right, none of those answers would be correct.

OpenStudy (anonymous):

i tyed it out right...

OpenStudy (anonymous):

i think it should be:\[\sin a (?)=\frac{1}{2}(\sin(a+b)+\sin(a-b))\]you have to figure out which goes in the question marks spot. If you expand the right side using the sum and difference formulas, it will be clear which is the right answer

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

right :)

OpenStudy (anonymous):

thanks :)